OpenStudy (anonymous):

what is the geometric mean of 1.5 and 12

OpenStudy (esshotwired):

To find the geometric mean of two numbers, it would be the square root of the product of the 2 numbers. This would be the formula to use: \[\sqrt{a _{1} \times a_{2}}\] a1 would be 1.5 and a2 would be 12 So solve: \[\sqrt{1.5\times12}\]

OpenStudy (anonymous):

Yes, and to extend this, for n values, if you have the product of n values , the geometric mean is the nth root of it.

OpenStudy (esshotwired):

1.5x12=18 so you would have to find the square root of 18. The way I do this is finding 2 factors of the number. So for this one, I would do:  Does this make sense?